OpenStudy (johnkim2330):

Determine how many, what type, and find the roots for f(x) = x3 - 5x2 – 25x + 125.

OpenStudy (johnkim2330):

@iPwnBunnies

OpenStudy (johnkim2330):

@OliverSt4rz

OpenStudy (praxer):

http://www.mathsisfun.com/algebra/fundamental-theorem-algebra.html this might help.

OpenStudy (johnkim2330):

i atill dont get it @praxer

OpenStudy (praxer):

$$\huge x^2(x−5)−25(x−5) = 0$$

OpenStudy (johnkim2330):

(x2-25)(x-5)?

OpenStudy (praxer):

see, for $$\huge x^3-5x^2$$ part, take $$\huge x^2 $$ common

OpenStudy (praxer):

than for $$\huge -25x+125$$ part take $$\huge -25$$ common.

OpenStudy (praxer):

then we will get $$\huge x^2(x-5)-25(x-5)=0$$ now take $$\huge (x-5) $$ common, than you get, $$ \huge (x^2-25) (x-5) =0 $$ finally you get, $$ \huge (x+5) (x-5)^2 $$ got it.

OpenStudy (praxer):

so, roots shall be $$\huge \pm 5$$